40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Wed May 25 16:56:44 2022, Paco, Configuration, BHD, IFO recovery - IMC alignment
|
Wed May 25 17:34:47 2022, yuta, Configuration, BHD, IFO recovery - IFO alignment
|
Thu May 26 19:55:43 2022, yuta, Configuration, BHD, Oplevs centered, BHD DCPDs are now online    
|
Fri May 27 17:45:53 2022, yuta, Configuration, BHD, BHD camera installed, GRY aligned
|
Thu Jun 2 20:05:37 2022, yuta, Configuration, PSL, IMC input power recovered to 1W, some alignment works 
|
Fri Jun 3 12:13:58 2022, Paco, Configuration, CDS, Fix RFM channels
|
Fri Jun 3 15:22:51 2022, yuta, Update, LSC, Both arms locked with POY/POX, IR beam centered on TMs with ASS 
|
Fri Jun 3 17:42:50 2022, yuta, Update, LSC, MICH locks with AS55_Q
|
Sun Jun 5 19:46:40 2022, Paco, Update, LSC, Fixed IMC Trans sum issue
|
Mon Jun 6 13:35:11 2022, Paco, Update, LSC, First calibrated spectra of MICH at AS55 Q 
|
Fri Jun 17 16:22:21 2022, yuta, Update, LSC, Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 6x
|
Thu Jul 7 18:18:19 2022, yuta, Update, LSC, Actuator calibration of ETMX and ETMX  
|
Thu Jul 7 18:22:12 2022, yuta, Update, LSC, Actuator calibration of MC2 using Yarm 
|
Fri Jul 8 16:18:35 2022, rana, Update, LSC, Actuator calibration of MC2 using Yarm
|
Mon Jul 18 17:07:12 2022, yuta, Update, LSC, x4.12 added to ETMX coil outputs to balance with ETMY
|
Thu Jun 16 18:23:15 2022, Paco, Configuration, BHD, Recovering LO beam in BHD DCPDs
|
Tue Jun 21 14:17:50 2022, yuta, Configuration, BHD, BHD DCPDs re-routed to c1sus2 
|
Tue Jul 19 16:00:34 2022, yuta, Configuration, BHD, Fast channels for BHD DCPDs now available in c1lsc but not in c1hpc  
|
Thu Jul 21 21:50:47 2022, Tega, Configuration, BHD, c1sus2 IPC update
|
Fri Jul 22 15:05:26 2022, Tega, Configuration, BHD, c1sus2 shared memory and ADC fix   
|
Fri Jul 22 17:46:10 2022, yuta, Configuration, BHD, c1sus2 watchdog update and DCPD ERR channels
|
Mon Jul 25 17:58:10 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update   
|
Mon Aug 1 18:42:39 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update  
|
Tue Aug 2 17:25:18 2022, Tega, Configuration, BHD, c1sus2 dolphin IPC issue solved 
|
Mon Jul 25 18:09:41 2022, Tega, Configuration, BHD, BHD Homodyne Phase control MEDM screen 
|
|
Message ID: 16877
Entry time: Thu May 26 19:55:43 2022
In reply to: 16875
Reply to this: 16880
16924
16932
|
Author: |
yuta |
Type: |
Configuration |
Category: |
BHD |
Subject: |
Oplevs centered, BHD DCPDs are now online |
|
|
[Paco, Yuta]
We have aligned the IFO (except for LO-AS and GRY), and centered all the oplevs.
We have also restored Gautam's in-air BHD DCPD setup and placed it to ITMY table.
BHD DC PD signals are now online at C1:XO4-MADC1_EPICS_CH4 and CH5.
Oplevs:
Aligned the IFO following the steps in elog 40m/16875.
When we were woking on BHD DCPDs, we lost REFL beam on camera and both arms flashing. Alignment was restored mostly with TT2 pitch.
We centered all the oplevs after the recovery (see attached).
BHD DCPDs:
1. We removed a circuit box with M2 ISS photodetector readout board from AP table, in-air BHD photodiodes from optics graveyard. (see LIGO-E2000436 and elog 40m/15493 for wiring diagram)
2. Taken out temporary two Thorlabs PDA100A used for aligning LO-AS during the vent from ITMY table, and placed the BHD setup in ITMY table (see attached and attached).
3. DB9 cable (15ft+10ft) was connected from M2 ISS box to anti-aliasing chassis for ADC1 of C1X04 at 1Y2 rack (see attached).
4. +/-18V power for M2 ISS box was supplied from 1Y1 rack.
5. BHD DCPD signals are now available at C1:XO4-MADC1_EPICS_CH4 and CH5 (see attached).
Next:
- Tweak alignment of green Y input to follow Yarm
- Do LO-AS alignment
- Centering of PDs everywhere with IFO aligned
- Update RTS model for BHD |
|
|
|
|
|